Tohban Report 2014-02-05

From RHESSI Wiki

Jump to: navigation, search

Tohban Reports
Start Date: 29 Jan 2014
End Date: 04 Feb 2014
Tohban: Juan C. Martinez Oliveros
Tohban email:
Next Tohban: Nicole Duncan
List all reports


Solar Activity

Solar activity was more than moderate this week. There are currently 5 ARs on the disk, the most active of which is NOAA 11967 which is responsible for the majority of activity this week. A major flare watch is in place, forecasting a flare higher than M5 in the following days.

How many GOES flares occurred?

 Flares above B, C, M, X class were      0    56    16     0

And how many of these are listed in the RHESSI flare list?

 Flares above B, C, M, X class were      0    38    11     0

And how many had EXCELLENT coverage?

 Flares above B, C, M, X class were      0    19     4     0

There were RHESSI flares/GOES flares 478 / 72 over the time range 29-Jan-14 05-Feb-14

Memory Management

The SSR was not emptying by the end of the daily pass set and reaching a maximum of ~49% and minimum of ~15%.

Spacecraft Status

The cold plate 1 and 2 are at about 116.3 K and 114.9 K respectively.

Data Gaps

No data gaps this week.

Detector issues

On 2014-Jan-29 at 22:11 UT, the G6 and G9 front fast thresholds were increased by 0x10 each.

  was     now

G6: 0x60 -> 0x70 G9: 0x30 -> 0x40

Detector 2 rear resets are been seen in the monitor rates with some sporadic blips in the rear slow valid rates.

Other notes

To mitigate the SSR fill level no front or rear events are been taken during eclipse.

Spacecraft Management

No operations this week

Decimation Active/Vigorous
HLAT Decimation Rear decimation weight 6, no front decimation
Night time data (fronts) Enabled 4 minutes before and after eclipse
Night time data (rears) Disabled, during entire eclipse
Require extra passes? No?
Requirement for moving pointer? No
Attenuator operation Normal
Detector problems? None
Personal tools